A rectangular poster has an area of 190 inches squared the height of a poster is one inch less than twice its width. Find in the dimensions of the poster.
Answer by JulietG(1812)

L * W = area
L * W = 190
L + 1 = 2W
Therefore, L = 2W - 1
The easiest way to do this is to find the factors, and then see what fits.
Factors of 190 = {190,1}, {95,2}, {5,38}, {19,10}
In which of these pairs is one number "twice as much less one" than the other?
.
.
.
.
The solution is 19,10. 19*10 = 190 square feet. The length (19) is twice the width (10) minus one.